Assume the markets for cherries, muffins and tea are each perfectly competitive. (a) In the market for cherries, the equilibrium price is $1 per kilogram and the equilibrium quantity is 1000 kilograms. Calculate the price elasticity of supply if price increases to $1.20 causing the quantity supplied to increase to 1500 kilograms. 


1
Expert's answer
2021-10-25T08:47:09-0400

% change in quantity supplied:

"=\frac{1500-1000}{1000}\times100=50"%

% change in price:

"=\frac{1.20-1}{1}\times100=20"%

Price Elasticity of Supply:

"=\frac { Percentage Change in Quantity Supplied}{ Percentage Change in Price}"

"=\frac{50}{20}"

"=2.5."